OpenStudy (nikkkk66666):

At which root does the graph of f(x) = (x + 4)6(x + 7)5 cross the x axis?

OpenStudy (sean_will21):

Need help?

OpenStudy (sean_will21):

im here to assist

OpenStudy (mathmale):

First of all, if you are presenting exponents, please do it correctly. Incorrect: f(x) = (x + 4)6(x + 7)5 correct: f(x) = (x + 4)^6(x + 7)^5 " ^ " represents "exponentiation."

OpenStudy (mathmale):

Actually, this function has two roots. What are they? Hint: Look up "root."